### v2

@@ Abstract
Developing large language models (LLMs) to cooperate and compete effectively within multi-agent systems (MASs) is a critical step towards more advanced intelligence. While reinforcement learning (RL) has proven effective for enhancing reasoning in single-agent tasks, its extension to multi-turn, multi-agent scenarios remains underexplored due to the challenges of long-horizon credit assignment and agent-specific advantage estimation. To address these challenges, we introduce MARSHAL, an end-to-end RL framework that incentivizes Multi-Agent Reasoning through Self-play witH strAtegic LLMs in both cooperative and competitive games. MARSHAL features a turn-level advantage estimator that aligns learning signals with each interaction for credit assignment, and an agent-specific advantage normalization to stabilize multi-agent training. By learning with self-play across cooperative and competitive games, MARSHAL agent trained from Qwen3-4B develops strong strategic abilities that generalize to held-out games with up to 28.728.7% performance improvements. More importantly, the capability acquired through self-play generalizes beyond games, yielding consistent performance gains of MASs in reasoning benchmarks. When integrated into leading MASs, our MARSHAL agent achieves significant performance gains of up to 10.010.0% on AIME, 6.66.6% on GPQA-Diamond, and 3.53.5% on average across all benchmarks. These results establish end-to-end RL training with self-play in strategic games as a powerful approach for developing generalizable multi-agent reasoning capabilities in LLMs.
@@ 1 Introduction
The remarkable capabilities of large language models (LLMs) have revolutionized numerous domains, enabling strong performance on a wide range of tasks from question answering to code generation (achiam2023gpt; team2023gemini). However, many real-world scenarios, such as negotiation (bianchi2024well), strategic gameplay (silver2016mastering; meta2022human), and collaborative software development (qian2023chatdev; wu2024autogen), inherently involve multiple agents interacting over long horizons. Enabling LLMs to cooperate and compete effectively within multi-agent systems (MASs) is a critical frontier for advancing artificial intelligence.
@@
While reinforcement learning (RL) has demonstrated remarkable success in enhancing the reasoning capabilities of individual LLMs (guo2025deepseek; team2025kimi), its extension to multi-turn, multi-agent tasks faces two critical challenges. First, the problem of long-horizon credit assignment arises in multi-turn interactions. As a sequence of actions can each result in a immediate reward while collectively leading to the final sparse reward, accurately attributing the contribution of the action in each turn is inherently more complex than standard single-turn settings. Second, multi-agent training couples heterogeneous game roles with asymmetric information and different payoff scales, which introduces variance in advantage estimation, destabilizing the training process.
@@
In this work, we present MARSHAL (Multi-Agent Reasoning through Self-play witH strAtegic LLMs), an end-to-end RL framework that incentivizes multi-agent reasoning skills via self-play in strategic games. We introduce two novel techniques to address the challenges in multi-turn, multi-agent self-play with Group-Relative Policy Optimization (GRPO) (shao2024deepseekmath). First, we propose a simple yet effective turn-level advantage estimator to enable fine-grained credit assignment. This allows the model to accurately attribute long-term outcomes to individual actions and provide learning signals across multiple turns and agents. Second, we propose an agent-specific advantage normalization that stabilizes the training process by calibrating advantage estimates relative to the performance of each agent. This normalization accounts for the heterogeneous roles in multi-agent systems and ensures stable policy updates. By integrating these components, MARSHAL enables robust self-play learning from game outcomes, empowering LLMs to develop strategic abilities and multi-agent reasoning skills through cooperation and competition with themselves.
@@
To evaluate the performance and generalization ability of MARSHAL agents, we conduct extensive experiments by training Qwen3-4B in a diverse range of cooperative and competitive games. Specifically, MARSHAL agents exhibit strong strategic ability across all game environments, with up to 28.728.7% performance improvements in three held-out games. More importantly, the capability acquired through self-play in games further generalizes to improvements of multi-agent systems in reasoning benchmarks. When integrated into both cooperative and competitive multi-agent systems, including AutoGen (wu2024autogen) and MAD (liang2023encouraging), our MARSHAL agents achieve performance improvements up to 10.010.0% on AIME, 6.66.6% on GPQA, and 3.53.5% on average across all benchmarks. We further conduct ablation studies to validate the effectiveness of our algorithmic design, complemented by a comprehensive analysis of reasoning patterns and failure modes to understand the successful generalization. The evaluation results establish MARSHAL as a powerful approach for developing generalizable multi-agent reasoning capabilities in LLMs.
@@
  • We propose MARSHAL, an end-to-end RL framework that enhances generalizable multi-agent reasoning capability through self-play in a diverse range of strategic games.
@@ 2 Preliminary
In standard RL fine-tuning for tasks like math solving, the environment is static. The model generates a complete Chain-of-Thought response, and only then is a terminal reward assigned. This process is typically modeled as a token-level Markov decision process (MDP). Here, an trajectory consists of a single "turn"—the generation of one complete response with multiple tokens. The state at step $t$, is the sequence of previously generated tokens $(o_{1},...,o_{t-1})$, and the action is the selection of the next token $o_{t}$. Given a task question $q$, the goal is to optimize the token-level policy $\pi(o_{t}|q,o_{<t})$ to produce a sequence that maximizes the final, sparse reward.
In contrast, strategic games introduce a more complex, hierarchical decision-making structure. An entire game, not a single response, constitutes an episode. This is best modeled as a turn-level MDP, where decisions occur at two levels. At the high level, the state $s_{k}$ represents the state of the game at the beginning of turn $k$ (e.g., the board configuration, cards in hand). A high-level action, $a_{k}$, corresponds to the agent’s entire output for that turn (e.g., "I place my X in the top-left corner and here is why…"). This action $a_{k}$ is itself a sequence of multiple tokens generated by the LLM’s low-level, autoregressive policy. In this case, the goal is to optimize the turn-level policy $\pi(a_{k}|s_{k})=\prod^{T}_{t=1}\pi(o_{k,t}|s_{k},o_{k,<t})$ to maximize the total return of all turns $R=\sum_{k=1}^{K}r_{k}$.
@@ 3 Method
In this section, we introduce MARSHAL, an end-to-end RL framework that enhances multi-agent reasoning ability of LLM through self-play in a diverse range of cooperative and competitive games. We begin by outlining the overall architecture of our training framework, building upon Group-Relative Policy Optimization (GRPO) (shao2024deepseekmath). We then detail our primary technical contributions for addressing the credit assignment and advantage estimation challenge in multi-turn, multi-agent training. Finally, we describe the selection of our game environments and design of reward structures. An overview of our proposed method is shown in Fig. 2.
@@ 3.1 Self-play with GRPO
To eliminate the extensive computational cost introduce by the critic model in classic Proximal Policy Optimization (PPO) (schulman2017proximal), GRPO rollout each query for multiple times, constructing a group of $G$ responses $\{o^{i}\}_{i=1}^{G}$ and their corresponding outcome rewards $\textbf{r}=\{r^{i}\}_{i=1}^{G}$, then estimate the advantage of each response by their relative reward to their corresponding group, which assigns equal advantage to the whole sequence (Fig.2 middle), yielding:
$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{q\sim P(Q),\{o^{i}\}_{i=1}^{G}\sim\pi_{\theta_{old}}(O|q)}\frac{1}{G}\sum^{G}_{i=1}\frac{1}{|o^{i}|}\sum^{|o^{i}|}_{t=1}\mathcal{J}_{\text{surr}}(\pi_{\theta};\pi_{\theta_{old}},A_{t}^{i},\varepsilon)$ (1)
@@
In multi-agent self-play, where all players in a strategic game is controlled by the same model, each episode of game results in a multi-turn trajectory for each player role, respectively. As a naive generalization of the original GRPO to our multi-turn scenario, we start by considering all trajectories from a game environment as a group of response, i.e. $\{(s_{k}^{i},a_{k}^{i})_{k=1}^{K^{i}}\}_{i=1}^{G}$, with total return as the terminal reward for each trajectory $\textbf{r}=\{R^{i}\}_{i=1}^{G}$. Then, with an additional summation for multiple turns, GRPO directly generalized to:
$\mathcal{J}_{\text{GRPO}}^{\text{multi}}(\theta)=\mathbb{E}_{s_{k}^{i}\sim P(S),o_{k,t}^{i}\sim\pi_{\theta_{old}}(O|s_{k}^{i},o_{k,<t}^{i})}\frac{1}{G}\sum^{G}_{i=1}\frac{1}{K^{i}}\sum^{K^{i}}_{k=1}\frac{1}{|o^{i}_{k}|}\sum^{|o^{i}_{k}|}_{t=1}\mathcal{J}_{\text{surr}}(\pi_{\theta};\pi_{\theta_{old}},A_{k,t}^{i},\varepsilon)$ (2)
@@ 3.2 Advantage estimation for multi-turn, multi-agent learning
While Group-Relative Policy Optimization (GRPO) has demonstrated remarkable efficiency and stability in single-turn settings, such direct application to the multi-turn, multi-agent structure of self-play introduces significant challenges of long-horizon credit assignment and agent-specific advantage estimation. To address these issues, we introduce two novel modifications to GRPO.
@@
To enable fine-grained credit assignment across a long trajectory, we incorporate the sequence of turn-level rewards $\mathbf{r}=\{(r_{k}^{i})_{k=1}^{K}\}_{i=1}^{G}$. This setup is analogous to the "Process Supervision" setting in the original GRPO paper (shao2024deepseekmath). In that context, the proposed advantage estimation involves first normalizing all rewards across the entire batch, $\tilde{r}_{k}^{i}=(r_{k}^{i}-\text{mean}(\mathbf{r}))/\text{std}(\mathbf{r})$, and then computing the cumulative sum of these normalized values, $A_{k}^{i}=\sum_{\hat{k}=k}^{K}\tilde{r}_{\hat{k}}^{i}$. While this procedure normalizes the immediate rewards, a critical flaw remains: the advantages $A_{k}^{i}$ are not, in general, zero-mean, which can reintroduce instability into training.
To rectify this, we propose a crucial inversion of these two steps: we first sum, then normalize (Fig.2 right). Specifically, we begin by calculating the standard Monte Carlo return (or cumulative reward) from each turn $k$ onwards: $R_{k}^{i}=\sum_{\hat{k}=k}^{K}r_{\hat{k}}^{i}$. We then compute the advantage by normalizing these returns to their mean: $A_{k,t}^{i}=R_{k}^{i}-\text{mean}(\textbf{R})$, where $\mathbf{R}$ is the collection of all cumulative rewards in the group. This formulation is equivalent to Generalized Advantage Estimation (GAE) (schulman2015high) with $\gamma=1$ and $\lambda=1$, where the state-value function $V(s_{k})$ is approximated by a simple, yet effective, baseline—the empirical mean of the batch returns $\mathbb{E}[\mathbf{R}]$. By ensuring the final advantages are properly centered, this method provides a much more stable learning signal for multi-turn decision-making.
@@
In many games, the expected return can be highly dependent on a player’s role (e.g., player 0 vs. player 1, or different roles in a cooperative game). Normalizing advantages across a heterogeneous batch of roles would force all players toward a shared baseline, which is statistically inappropriate and can wash out role-specific learning signals.
To address this, we refine our advantage calculation further. We partition the batch of trajectories into sub-groups based on player role and apply the turn-level advantage estimator described above independently within each sub-group $G^{p}$ (Fig.2 right), where $p$ denotes the player index. This ensures that the advantage for a given action is calculated relative to the average outcome for that specific role, providing a more accurate and stable credit assignment in multi-agent settings. This agent-specific, sum-then-normalize advantage estimation forms the final objective for MARSHAL:
@@
$\frac{1}{P}\sum^{P}_{p=1}\frac{1}{G_{p}}\sum^{G_{p}}_{i=1}\frac{1}{K^{i}}\sum^{K^{i}}_{k=1}\frac{1}{|o^{p,i}_{k}|}\sum^{|o^{p,i}_{k}|}_{t=1}\mathcal{J}_{\text{surr}}(\pi_{\theta};\pi_{\theta_{old}},A_{k,t}^{p,i},\varepsilon)$
@@ 3.3 Game environments
To cultivate complementary multi-agent skills for MARSHAL, we design a portfolio of six strategic, two-player games. These games are partitioned into a training set and a more complex, held-out testing set for out-of-distribution (OOD) evaluation:
  • Perfect-information, competitive games: To build deterministic planning and role adaptation, we train on Tic-Tac-Toe, requiring the agent to recognize its strategic position (e.g., first-mover vs. second-mover) and plan accordingly. For evaluation, Connect Four serves as a more complex out-of-distribution test due to its vastly larger state space.
  • Imperfect-information, competitive games: To foster robust reasoning and decision-making under imperfect-information and uncertainty, we train on Kuhn Poker, a simplified poker variant. We evaluate generalization on the more sophisticated Leduc Hold’em.
  • Imperfect-information, cooperative games: The inclusion of cooperative games is a key differentiator of our work. We use Hanabi to develop social intelligence skills like intent recognition and Theory of Mind. Specifically, we train on Mini Hanabi and evaluate on a more challenging variant (Simple Hanabi) to test the generalization of cooperative strategies.
Collectively, this diverse portfolio of both competitive and cooperative challenges ensures the agent develops a robust multi-agent skillset and avoids overfitting to the narrow, competitive-only mindset of prior work (liu2025spiral).

### v3

@@ Abstract
Developing Large Language Models (LLMs) to cooperate and compete effectively within multi-agent systems (MASs) is a critical step towards more advanced intelligence. While reinforcement learning (RL) has proven effective for enhancing reasoning in single-agent tasks, its extension to multi-turn, multi-agent scenarios remains underexplored due to the challenges of long-horizon credit assignment and agent-specific advantage estimation. To address these challenges, we introduce MARSHAL, an end-to-end RL framework that incentivizes Multi-Agent Reasoning through Self-play witH strAtegic LLMs in both cooperative and competitive games. MARSHAL features a turn-level advantage estimator that aligns learning signals with each interaction for credit assignment, and an agent-specific advantage normalization to stabilize multi-agent training. By learning with self-play across cooperative and competitive games, MARSHAL agents trained from Qwen3-4B develop strong strategic abilities, with up to 28.728.7% performance improvements in held-out games. More importantly, the capability acquired through self-play generalizes beyond games, yielding consistent performance gains of MASs in reasoning benchmarks. When integrated into leading MASs, our MARSHAL agent achieves significant zero-shot performance gains of up to 10.010.0% on AIME, 7.67.6% on GPQA-Diamond, and 3.53.5% on average across all benchmarks. These results establish self-play in strategic games as a powerful approach for developing generalizable multi-agent reasoning capabilities in LLMs.
@@ 1 Introduction
The remarkable capabilities of Large Language Models (LLMs) have revolutionized numerous domains, enabling strong performance on a wide range of tasks from question answering to code generation (Achiam et al., 2023; Team et al., 2023). However, many real-world scenarios, such as negotiation (Bianchi et al., 2024), strategic gameplay (Silver et al., 2016; FAIR et al., 2022), and collaborative software development (Qian et al., 2023; Wu et al., 2024a) inherently involve multiple agents interacting over long horizons. Enabling LLMs to cooperate and compete effectively within multi-agent systems (MASs) is a critical frontier for advancing artificial intelligence.
@@
While reinforcement learning (RL) has demonstrated remarkable success in enhancing the reasoning capabilities of individual LLMs (Guo et al., 2025; Team et al., 2025), its extension to multi-turn, multi-agent tasks faces two critical challenges. First, the problem of long-horizon credit assignment arises in multi-turn interactions. As a sequence of actions can each result in an immediate reward while collectively leading to the final sparse reward, accurately attributing the contribution of the action in each turn is inherently more complex than standard single-turn settings. Second, multi-agent training couples heterogeneous game roles with asymmetric information and different payoff scales, which introduces variance in advantage estimation, destabilizing the training process.
@@
In this work, we present MARSHAL (Multi-Agent Reasoning through Self-play witH strAtegic LLMs), an end-to-end RL framework that incentivizes multi-agent reasoning capabilities via self-play in strategic games. We introduce two novel techniques to address the challenges in multi-turn, multi-agent self-play with Group-Relative Policy Optimization (GRPO) (Shao et al., 2024). First, we propose a simple yet effective turn-level advantage estimator to enable fine-grained credit assignment. This allows the model to accurately attribute long-term outcomes to individual actions and provide learning signals across multiple turns and agents. Second, we propose an agent-specific advantage normalization that stabilizes the training process by calibrating advantage estimates relative to the performance of each agent. This normalization accounts for the heterogeneous roles in multi-agent systems and ensures stable policy updates. By integrating these components, MARSHAL enables robust self-play learning from game outcomes, empowering LLMs to develop strategic abilities and multi-agent reasoning skills through cooperation and competition with themselves.
@@
To evaluate the performance and generalization ability of MARSHAL agents, we conduct extensive experiments by training Qwen3-4B in a diverse range of cooperative and competitive games. Specifically, MARSHAL agents exhibit strong strategic ability across all game environments, with up to 28.728.7% performance improvements in three held-out games. More importantly, the capability acquired through self-play in games further generalizes to improvements of multi-agent systems in reasoning benchmarks. When integrated into both cooperative and competitive multi-agent systems, including AutoGen (Wu et al., 2024a) and MAD (Liang et al., 2023), our MARSHAL agents achieve zero-shot performance improvements of up to 10.010.0% on AIME, 7.67.6% on GPQA, and 3.53.5% on average across all benchmarks. We further conduct ablation studies to validate the effectiveness of our algorithmic design, complemented by a comprehensive analysis of reasoning patterns and failure modes to understand the successful generalization. The evaluation results establish MARSHAL as a powerful approach for developing generalizable multi-agent reasoning capabilities in LLMs.
@@
  • We propose MARSHAL, an end-to-end RL framework that enhances multi-agent reasoning through self-play in a diverse range of strategic games.
@@ 2 Preliminary
In standard RL fine-tuning for tasks like math, the environment is static. The model generates a complete Chain-of-Thought response, and only then is a terminal reward assigned. This process is typically modeled as a token-level Markov decision process (MDP). Here, a trajectory consists of a single "turn", which is the generation of one complete response with multiple tokens. The state at step $t$ is the sequence of previously generated tokens $(o_{1},...,o_{t-1})$, and the action is the selection of the next token $o_{t}$. Given a task question $q$, the goal is to optimize the token-level policy $\pi(o_{t}|q,o_{<t})$ to produce a sequence that maximizes the final sparse reward.
In contrast, strategic games introduce a more complex, hierarchical decision-making structure. An entire game, not a single response, constitutes an episode. This is best modeled as a turn-level MDP, where decisions occur at two levels. At the high level, the state $s_{k}$ represents the state of the game at the beginning of turn $k$ (e.g., the board configuration, the cards in hand, etc.). A high-level action, $a_{k}$, corresponds to the agent’s entire output for that turn (e.g., "I place my X in the top-left corner and here is why…"). This action $a_{k}$ is itself a sequence of multiple tokens generated by the LLM’s low-level, autoregressive policy. In this case, the goal is to optimize the turn-level policy $\pi(a_{k}|s_{k})=\prod^{T}_{t=1}\pi(o_{k,t}|s_{k},o_{k,<t})$ to maximize the total return of all turns $R=\sum_{k=1}^{K}r_{k}$.
@@ 3 Method
In this section, we introduce MARSHAL, an end-to-end RL framework that enhances multi-agent reasoning ability of LLM through self-play in a diverse range of cooperative and competitive games. We begin by outlining the overall architecture of our training framework, building upon Group-Relative Policy Optimization (GRPO) (Shao et al., 2024). We then detail our primary technical contributions for addressing the credit assignment and advantage estimation challenge in multi-turn, multi-agent training. Finally, we describe the selection of our game environments and design of reward structures. An overview of our proposed method is shown in Fig. 2.
@@ 3.1 Self-play with GRPO
To eliminate the extensive computational cost introduce by the critic model in classic Proximal Policy Optimization (PPO) (Schulman et al., 2017), GRPO rollout each query for multiple times, constructing a group of $G$ responses $\{o^{i}\}_{i=1}^{G}$ and their corresponding outcome rewards $\textbf{r}=\{r^{i}\}_{i=1}^{G}$, then estimate the advantage of each response by their relative reward to their corresponding group, which assigns equal advantage to the whole sequence (Fig.2 middle), yielding
$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{q\sim P(Q),\{o^{i}\}_{i=1}^{G}\sim\pi_{\theta_{old}}(O|q)}\frac{1}{G}\sum^{G}_{i=1}\frac{1}{|o^{i}|}\sum^{|o^{i}|}_{t=1}\mathcal{J}_{\text{surr}}(\pi_{\theta};\pi_{\theta_{old}},A_{t}^{i},\varepsilon),$ (1)
@@
In multi-agent self-play, where all players in a strategic game is controlled by the same model, each episode of game results in a multi-turn trajectory for each player role, respectively. As a naive generalization of the original GRPO to our multi-turn scenario, we start by considering all trajectories from a game environment as a group of response, i.e. $\{(s_{k}^{i},a_{k}^{i})_{k=1}^{K^{i}}\}_{i=1}^{G}$, with total return as the terminal reward for each trajectory $\textbf{r}=\{R^{i}\}_{i=1}^{G}$. Then, with an additional summation for multiple turns, GRPO directly generalized to
$\mathcal{J}_{\text{GRPO}}^{\text{multi}}(\theta)=\mathbb{E}_{s_{k}^{i}\sim P(S),o_{k,t}^{i}\sim\pi_{\theta_{old}}(O|s_{k}^{i},o_{k,<t}^{i})}\frac{1}{G}\sum^{G}_{i=1}\frac{1}{K^{i}}\sum^{K^{i}}_{k=1}\frac{1}{|o^{i}_{k}|}\sum^{|o^{i}_{k}|}_{t=1}\mathcal{J}_{\text{surr}}(\pi_{\theta};\pi_{\theta_{old}},A_{k,t}^{i},\varepsilon),$ (2)
@@ 3.2 Advantage estimation for multi-turn, multi-agent learning
While GRPO has demonstrated remarkable efficiency and stability in single-turn settings, such direct application to the multi-turn, multi-agent structure of self-play introduces significant challenges of long-horizon credit assignment and agent-specific advantage estimation. To address these issues, we introduce two novel modifications to GRPO.
@@
To enable fine-grained credit assignment across a long trajectory, we incorporate the sequence of turn-level rewards $\mathbf{r}=\{(r_{k}^{i})_{k=1}^{K}\}_{i=1}^{G}$. This setup is analogous to the "Process Supervision" setting in the original GRPO paper (Shao et al., 2024). In that setting, the proposed advantage estimation involves first normalizing all rewards across the entire batch, $\tilde{r}_{k}^{i}=(r_{k}^{i}-\text{mean}(\mathbf{r}))/\text{std}(\mathbf{r})$, and then computing the cumulative sum of these normalized values, $A_{k}^{i}=\sum_{\hat{k}=k}^{K}\tilde{r}_{\hat{k}}^{i}$. However, since the distributions of intermediate rewards can vary significantly, treating the entire set of rewards as a single distribution for global normalization is likely inappropriate and potentially problematic.
To address this problem, we propose a crucial inversion of these two steps: we first sum, then normalize (Fig.2 right). Specifically, we begin by calculating the standard Monte Carlo return (or cumulative reward) from each turn $k$ onwards: $R_{k}^{i}=\sum_{\hat{k}=k}^{K}r_{\hat{k}}^{i}$. We then compute the advantage by normalizing these returns to their mean: $A_{k,t}^{i}=R_{k}^{i}-\text{mean}(\textbf{R})$, where $\mathbf{R}$ is the collection of all cumulative rewards in the group. This formulation is equivalent to Generalized Advantage Estimation (GAE) (Schulman et al., 2015) with $\gamma=1$ and $\lambda=1$, where the value function $V(s_{k})$ is approximated by a simple yet effective baseline: the empirical mean of the batch returns $\mathbb{E}[\mathbf{R}]$. By ensuring the final advantages are properly centered, this method provides a much more stable learning signal for multi-turn decision-making.
@@
In many games, the expected return can be highly dependent on a player’s role (e.g., player 0 vs. player 1, or different roles in a cooperative game). Normalizing advantages across different roles would force all players toward a shared baseline, which is statistically inappropriate and can obscure role-specific learning signals.
To address this, we refine our advantage calculation further. We partition the batch of trajectories into sub-groups based on player role and apply the turn-level advantage estimator described above independently within each sub-group $G^{p}$ (Fig.2 right), where $p$ denotes the player index. This ensures that the advantage for a given action is calculated relative to the average outcome for that specific role, providing a more accurate and stable credit assignment in multi-agent settings. This agent-specific, sum-then-normalize advantage estimation forms the final objective for MARSHAL as
@@
$\frac{1}{P}\sum^{P}_{p=1}\frac{1}{G_{p}}\sum^{G_{p}}_{i=1}\frac{1}{K^{i}}\sum^{K^{i}}_{k=1}\frac{1}{|o^{p,i}_{k}|}\sum^{|o^{p,i}_{k}|}_{t=1}\mathcal{J}_{\text{surr}}(\pi_{\theta};\pi_{\theta_{old}},A_{k,t}^{p,i},\varepsilon),$
@@ 3.3 Game environments
To incentivize comprehensive multi-agent reasoning ability for MARSHAL, we select a range of six strategic, two-player games. These games are partitioned into a training set and a more complex, held-out testing set for out-of-distribution (OOD) evaluation:
  • Perfect-information, competitive games: To enable deterministic planning and role adaptation, we train on Tic-Tac-Toe, requiring the agent to recognize its strategic position (e.g., first-mover vs. second-mover) and plan accordingly. For evaluation, Connect Four serves as a more complex out-of-distribution test due to its vastly larger state space.
  • Imperfect-information, competitive games: To foster robust reasoning and decision-making under imperfect information and uncertainty, we train on Kuhn Poker, a simplified poker variant. We evaluate generalization on the more sophisticated Leduc Hold’em.
  • Imperfect-information, cooperative games: To develop social intelligence like intent recognition and Theory of Mind, we consider the classic cooperative card game Hanabi. Specifically, we train on Mini Hanabi and evaluate on a more challenging variant Simple Hanabi to test the generalization of cooperative strategies.
Collectively, this diverse range of both competitive and cooperative games ensures the LLM agent develops generalizable multi-agent reasoning capabilities.